# Human-Agent Joint Learning for Efficient Robot Manipulation Skill Acquisition

### v2

@@ 4 Experiments
Table 1: User studies on three dexterous hand tasks.
@@
In Tab. 1, by sharing control between humans and learned agents, our system shows improvements in both success rate and collection speed, while the average horizon length of the collected trajectories is reduced. This suggests that our system enhances the efficiency of data collection by facilitating a process that is easier to succeed, faster, and more fluid in terms of trajectory smoothness. To ensure the fairness of the experiment wasn’t compromised, we equally divided the user group into two parts, Group 1 first collected data directly by themselves (w/o Ours) and then collected data with an assistive agent (w/ Ours), while the Group 2 reversed the order, first (w/ Ours) mode and then (w/o Ours) mode.
@@
To gain deeper insight into how the learned agent assists the human operator, we visualize several keyframes from the data collection process of three dexterous hand tasks under shared control mode with $\gamma=0.4$. From Fig. 9, it is evident that human operators are not required to provide too precise control with the assistive agent facilitating shared control over the dexterous hand. Instead, they only need to convey high-level intentions, such as the direction of hand movement or finger grasp motions. In multi-stage tasks, like picking up a hammer and then using it to drive a nail, operators only need to provide a trigger action to guide the agent to transition from one sub-stage to the next. As a result, less effort and attention are required, making the data collection easier to execute successfully and speeding it up.
@@
Quantitative Evaluation. When the learned agent shares control with users, the system effectively corrects imperfect human control signals to achieve specific tasks. Given the challenge of directly measuring the level of imperfection in user signals and the correction ability of our system, we simulate the human using a baseline agent trained with Behavior Cloning (BC). Additionally, we import noise to the agent’s control signal through the diffusion policy’s forward diffuse process.
@@
In Fig. LABEL:fig:simulateduser, the x-axis represents the forward ratio, where a higher forward ratio corresponds to the lower quality of action provided by the simulated operator and a higher control ratio of the learned agent on the dexterous hand. The graph illustrates that with limited data availability, the agent can assist the simulated operator more effectively. As the agent accesses and trains with more data, its ability to correct actions improves. The results show that our system gradually diminishes the demand for the operator’s attention and effort, thereby enhancing the efficiency of data collection.
@@
Furthermore, once sufficient data is collected and the assistive agent is trained, it can transition into full autonomy mode by setting $\gamma$ to 1.0 and denoising random actions from the Gaussian distribution noise. Across three different dexterous manipulation tasks, we can achieve success rates of 0.76, 0.78, and 0.89, indicating that the assistive agent can effectively transform into an automated dexterous manipulation agent.
@@
From our experiments, we have observed that the assistive agent significantly aids human operators in managing fine control at the low level, especially in scenarios where accurate observation by humans is challenging, complicating effective action control. For instance, tasks such as grasping an egg or moving a hammer present visual challenges. It can be difficult to visually confirm whether the egg is securely grasped or if there’s a risk of it being dropped. This uncertainty makes it hard for human operators to react promptly to sudden changes. However, within our proposed joint learning framework, human operators are primarily required to focus on high-level intentions and task planning during data collection, while the assistive agent manages the detailed low-level actions. This division of labor significantly reduces the burden on human operators by clearly separating strategic planning from execution tasks, streamlining the collaboration between humans and machines.
@@
Table 2: Data quality on downstream tasks.
@@
Table 3: Tool-Use and Push-cube task success rate under increasing data.
Data Quality on Downstream Task. In this section, we illustrate that collecting data under shared control does not compromise the quality of the data. We gather dexterous hand and gripper manipulation demonstrations via the proposed system in two modes: fully controlling the robots by a human ($\mathcal{H}$) and sharing control ($\mathcal{S}$) between the human operator and the learned assistive agent. And utilize these data to train different kinds of agents, like BC, BC-RNN [7], and Diffusion Policy (DP) [8].
In Tab. 2, compared to directly collecting human demonstrations from the expert human operator, who can achieve success rates and efficiency comparable to those with agent assistance, the data collected by sharing control between the human and the assistive agent can achieve comparable or even surprisingly better results with BC and BC-RNN. Their results are comparable with DP, possibly as DP can better fit the tasks, which is in line with [8].
In Tab. 3, we compare the effects of using different sets of data to train BC and DP. We can find that utilizing more data collected under the shared control mode leads to comparable performance on the tool-use task. This verifies that the new data contributes significantly to policy learning and can achieve a similar effect compared to the data from human experts but at a much lower cost. These results indicate that the data collected under the proposed paradigm have sufficient quality and efficiency for downstream tasks.
Figure 12: Real world setting. 1. Pick-and-Place: (a) Hardware setup. The robot gripper should pick the red pot up (b-c) and place it onto the black induction cooker (d). 2. Articulated-Manipulation: use a gripper to open the drawer. 3. Push-cube: use a gripper to push the cube across the line.
@@
Table 4: Real world parallel gripper Pick-and-Place task user study.
@@
Table 5: Real world parallel gripper experiments of data quality.
Real World Experiment. To better evaluate our system, we further conduct real-world experiments. Three tasks are adopted: Pick-and-Place, Articulated-Manipulation, and Push-cube in Fig. 15. Following the same rules as Sec. 4, four human volunteers are invited to participate in the user study to collect data under two modes: one where control is shared between the human operator and the learned agent (w/ Ours), and the other where control is directly by the human operator alone (w/o Ours). Our proposed system achieves significant improvements in success rate and collection speed by sharing control between human operators and learned agents, as demonstrated in Tab. 4. Additionally, data gathered under our proposed joint learning shared control mode yield performance on the three tasks that are comparable to those pure human datasets using BC and DP, further substantiated by the results presented in Tab. 5.
5 Conclusion
@@
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grants 62306175.

### v3

@@ IV EXPERIMENTS
TABLE I: User studies on three dexterous hand tasks.
@@
IV-B Efficiency of Data Collection.
Our proposed system leverages shared control between human operators and learned agents to enhance the efficiency of data collection. To learn how the assistant agent could improve the data collection process, we conducted a user study.
In the user study, 10 human operators participate, collecting data under two modes: one where control is shared between the operator and the learned agent (w/ Ours), and the other where control is directly by the operator alone (w/o Ours). Each participant is instructed to collect as much data as possible within three minutes under two different modes for three dexterous hand tasks. Three metrics are evaluated: Success Rate (Percent) indicates the percentage of attempts where data collection was successful. Horizon Length (Steps per Sample) measures the length of each collected trajectory, with a lower horizon length indicating smoother data collection. Collection Speed (Samples per Hour) refers to the number of successful trajectories that can be collected in one hour.
@@
In Tab. I, by sharing control between humans and learned agents, our system shows improvements in both success rate and collection speed, while the average horizon length of the collected trajectories is reduced. This suggests that our system enhances the efficiency of data collection by facilitating a process that is easier to succeed, faster, and more fluid in terms of trajectory smoothness. To ensure the fairness of the experiment wasn’t compromised, we equally divided the user group into two parts, Group 1 first collected data directly by themselves (w/o Ours) and then collected data with an assistive agent (w/ Ours), while the Group 2 reversed the order, first (w/ Ours) mode and then (w/o Ours) mode.
IV-C Quantitative Evaluation.
@@
To gain deeper insight into how the learned agent assists the human operator, we visualize several keyframes from the data collection process of three dexterous hand tasks. From Fig. 11, it is evident that human operators are not required to provide too precise control with the assistive agent facilitating shared control over the dexterous hand. Instead, they only need to convey high-level intentions, such as the direction of hand movement or finger grasp motions. In multi-stage tasks, like picking up a hammer and then using it to drive a nail, operators only need to provide a trigger action to guide the agent to transition from one sub-stage to the next. As a result, less effort and attention are required, making the data collection easier to execute successfully and speeding it up.
@@
When the learned agent shares control with users, the system effectively corrects imperfect human control signals to accomplish specific tasks. Given the challenge of directly measuring the level of imperfection in user signals and the correction ability of our system, we simulate human input using a baseline agent trained with Behavior Cloning (BC) as a proxy for user control.
@@
In Fig. LABEL:fig:simulateduser, additional data collected under our framework effectively contributes to training improvements. The graph illustrates that with limited data availability, the agent can assist the simulated operator more effectively. As the agent gains access to and trains on more data, its ability to correct actions improves. These results indicate that our system gradually reduces the demand for the operator’s attention and effort, thereby enhancing the overall efficiency of data collection process.
@@
Furthermore, once sufficient data is collected and the assistive agent is trained, it can transition into full autonomy mode by setting $\gamma=1$ and denoising actions from the noise sampled from Gaussian distributions. Across three different dexterous manipulation tasks, we can achieve success rates of 0.76, 0.78, and 0.89, indicating that the assistive agent can effectively transform into an automated dexterous manipulation agent.
@@
From our experiments, we have observed that the assistive agent significantly aids human operators in managing fine control, especially in scenarios where accurate observation by humans is challenging. For instance, tasks such as grasping an egg or moving a hammer present visual challenges. It can be difficult to visually confirm whether the egg is securely grasped or if there’s a risk of it being dropped. This uncertainty makes it hard for human operators to react promptly to sudden changes. However, within our proposed joint learning framework, human operators are primarily required to focus on high-level intentions and task planning during data collection, while the assistive agent manages the detailed low-level actions. This division of labor significantly reduces the burden on human operators by clearly separating strategic planning from execution tasks, streamlining the collaboration between humans and machines.
@@
TABLE II: Data quality on downstream tasks.
@@
TABLE III: Agent performance under increasing data.
IV-D Data Quality on Downstream Task.
In this section, we illustrate that collecting data under shared control does not compromise the quality of the data. We gather dexterous hand and gripper manipulation demonstrations via the proposed system in two modes: fully controlling the robots by a human ($\mathcal{H}$) and sharing control ($\mathcal{S}$) between the human operator and the learned assistive agent. And utilize these data to train different kinds of agents, like BC, BC-RNN [7], and Diffusion Policy (DP) [8].
In Tab. II, compared to directly collecting human demonstrations from the expert human operator, who can achieve success rates and efficiency comparable to those with agent assistance, the data collected by sharing control between the human and the assistive agent can achieve comparable or even surprisingly better results with BC and BC-RNN. Their results are comparable with DP, possibly as DP can better fit the tasks, which is in line with [8].
In Tab. III, we compare the effects of using different sets of data to train BC and DP. We can find that utilizing more data collected under the shared control mode leads to comparable performance on the tool-use and push-cube tasks. This verifies that the new data contributes significantly to policy learning and can achieve a similar effect compared to the data from human experts but at a much lower cost. These results indicate that the data collected under the proposed paradigm have sufficient quality and efficiency for downstream tasks.
IV-E Real World Experiment and User Feedback.
To better evaluate our system, we further conduct real-world experiments. Three tasks are adopted: Pick-and-Place, Articulated-Manipulation, and Push-cube in Fig. 14. Following the same rules as Sec. IV-B, four human volunteers are invited to participate in the user study to collect data under two modes: one where control is shared between the human operator and the learned agent (w/ Ours), and the other where control is directly by the human operator alone (w/o Ours). Our proposed system achieves significant improvements in success rate and collection speed by sharing control between human operators and learned agents, as demonstrated in Tab. IV. Additionally, data gathered under our proposed joint learning shared control mode yield performance on the three tasks that are comparable to those pure human datasets using BC and DP, further substantiated by the results presented in Tab. V.
Figure 14: Real world setting. 1. Pick-and-Place: use the gripper to pick the red pot up and place it onto the black induction cooker. 2. Articulated-Manipulation: use the gripper to open the drawer. 3. Push-cube: use the gripper to push the cube across the black line.
@@
TABLE IV: Real world gripper Pick-and-Place task user study.
@@
TABLE V: Real world gripper experiments of data quality.
Satisfaction: $\alpha=0.769$
1. It is fun to use.
2. It works the way I want it to work.
3. It is wonderful.
4. It helps me be more effective.
5. It is flexible.
User-Friendly: $\alpha=0.852$
6. It is simple to use.
7. It is effortless.
8. I can use it without written instructions.
9. I do not notice any inconsistencies as I use it.
TABLE VI: Subjective Measures.
We have developed a questionnaire comprising shown in Tab. VI to capture various dimensions of user experience and ergonomics, and we invited 10 volunteers to rate our system based on their feedback.
This questionnaire assesses ease of use and overall satisfaction. The reliability of our questionnaire is supported by strong Cronbach’s alpha values: $\alpha=0.769$ for the satisfaction section and $\alpha=0.852$ for the user-friendly section, indicating internal consistency.
V CONCLUSION